# Improving Multilingual Capabilities with Cultural and Local Knowledge in Large Language Models While Enhancing Native Performance

### v2

@@ 9.2 Model biases over choices
The observations from domain wise performance changes by Phi and Qwen were significantly different. The domains which were well represented in our training data had a significant boost on both languages of MMLU. Despite training on MCQs which consist of 2-4 options, similar results of improvement were seen over MMLU-Pro which has upto 10 options. On the other hand, Phi-4 had a higher performance boost over MMLU which has the same number of options as the samples in the training data, but the performance over MMLU-Pro dropped irrespective of domain. The distribution of choices made by each of our LLMs and the corresponding original implementation can be seen in Figure 2. The instruction tuning dataset we used had an equal distribution of each of the choices among MCQ samples. The original Qwen model overwhelmingly chose from the final two options while our model was able to generalize well despite not being trained on MCQs with 10 choices. On the other hand the original phi-4 was able to perform better than its counterpart, but despite being fine-tuned with equal distribution of choices, the model displayed an inclination towards the first choice among the list of options. The extent of this bias varied between each domain significantly. More on this can be seen in Appendix D. As our phi model was fine-tuned from the original models’ instruct variants, the biases were assumed to have been carried forward. Our models were able to respond well with less biases in choices over the domains whose samples are present in large quantities in our training data. To further look into this, we tried to fine-tune the base variant of qwen-2.5-14B rather than the instruct model to see the choice distribution over MMLU-Pro, while most of our dataset’s samples of MCQs were having 4-5 samples, it was reflected in the choices made as seen in Figure 3 which demonstrates the issue within the original model similar to previous works demonstrating sensitivity on models’ sensitivity to order of choices (Pezeshkpour and Hruschka, 2024). But a well balanced instruction tuning dataset can minimize this issue or an evaluation independent of order of choices (Zheng et al., 2023). A slight tilt from left to right in Figure 2 and Figure 3 can be expected as not all questions are accompanied by 10 options with a considerable amount having less.
@@ 10.1 Scalability to other languages
As not every language has readily available datasets of even a few domains, we took an approach of using just translated datasets for all domains other than those used for localized and cultural knowledge addition. This would enable reusing the approach to build bi-lingual LLMs optimized for other languages as long as a proficient LLM supports the language to translate the texts fluently.
@@ 10.2 Model Efficiency
Unsloth’s version of phi-4 (Unsloth AI, 2023) with llama architecture led to an improved performance but increased emissions. Our model resulted in lesser emissions during evaluation over the open-llm-leaderboard, while improving the model’s performance. A comparison of our model to the original and unsloth’s phi-4 can be seen in Figure 4.
11 License
Our Qwen and Phi models are available through the same licenses as the models we used as a base i.e apache-2.0 and mit respectively. the models can be accessed here  [^3] Our Phi-4 model :https://huggingface.co/. The training datasets are publicly available here [^4] Datasets : https://huggingface.co/. Most datasets used for training the models have a copyleft license, with the rest having no license specified and are publicly available on huggingface.
@@ 11 License
Acknowledgments
This work was supported by a prior research grant from Traversaal.ai
@@ Appendix A Model Replication
The hyper-parameters used for training can be seen below in Table 11. The initial training attempts using a portion of the data (i.e 8% samples) were done on various different devices, the final models were trained on a single H200 SXM for 55,56,54 hours each using Qwen2.5-14B-Instruct, Phi-4, Qwen2.5-14B-base respectively.
@@
Figure 4: Emissions : open-llm-leaderboard evaluation
Hyperparameter Value
@@
Domain Dataset Total Used Hindi Original
Samples Samples Ratio Source
@@
The initially collected dataset sources, sample sizes and the later used sample counts can be seen in Table 12 along with the ratios of each language. The sampling within each dataset is done at random using the seed specified in Table 11. The samples were sorted in ascending order based on input size and the longest 600 samples in terms of input token count were added in the beginning of the training data.
Appendix B Datasets and Benchmarks Info
@@
Appendix C Results from other attempts
@@
Benchmarks Ratio of ARC-Challenge ARC-Easy MMLU BoolQ Context-MCQ Overall Average
Data used? Hindi En Hi En Hi En Hi En Hi En Hi En Hi Tot
@@
Benchmarks Ratio of ARC-Challenge ARC-Easy MMLU BoolQ Context-MCQ Overall Average
Data used? Hindi En Hi En Hi En Hi En Hi En Hi En Hi Tot
@@
Benchmarks Ratio of ARC-Challenge ARC-Easy MMLU BoolQ Context-MCQ Overall Average
Data used? Hindi En Hi En Hi En Hi En Hi En Hi En Hi Tot
@@
Benchmarks Ratio of ARC-Challenge ARC-Easy MMLU BoolQ Context-MCQ Overall Average
Data used? Hindi En Hi En Hi En Hi En Hi En Hi En Hi Tot
@@
Benchmarks Ratio of ARC-Challenge ARC-Easy MMLU BoolQ Context-MCQ Overall Average
Data used? Hindi En Hi En Hi En Hi En Hi En Hi En Hi Tot
@@
Appendix D Model Choices

### v3

@@ 9.2 Model biases over choices
The observations from domain-wise performance changes by Phi and Qwen were significantly different. The domains that were well represented in our training data had a significant boost in both languages of MMLU. Despite training on MCQs, which consist of 2-4 options, similar results of improvement were seen over MMLU-Pro, which has up to 10 options. On the other hand, Phi-4 had a higher performance boost over MMLU, which has the same number of options as the samples in the training data, but the performance over MMLU-Pro dropped irrespective of domain. The distribution of choices made by each of our LLMs and the corresponding original implementation can be seen in Figure 2. The instruction tuning dataset we used had an equal distribution of each of the choices among MCQ samples. The original Qwen model overwhelmingly chose from the final two options, while our model was able to generalize well despite not being trained on MCQs with 10 choices. On the other hand, the original phi-4 was able to perform better than its counterpart, but despite being fine-tuned with equal distribution of choices, the model displayed an inclination towards the first choice among the list of options. The extent of this bias varied between each domain significantly. More on this can be seen in Appendix E. As our Phi model was fine-tuned from the original models’ instruct variants, the biases were assumed to have been carried forward. Our models were able to respond well with fewer biases in choices over the domains whose samples are present in large quantities in our training data. To further look into this, we tried to fine-tune the base variant of qwen-2.5-14B rather than the instruct model to see the choice distribution over MMLU-Pro. While most of our dataset’s samples of MCQs had 4-5 samples, it was reflected in the choices made as seen in Figure 4, which demonstrates the issue within the original model similar to previous works demonstrating sensitivity on models’ sensitivity to order of choices (Pezeshkpour and Hruschka, 2024). But a well-balanced instruction-tuning dataset can minimize this issue or an evaluation independent of the order of choices (Zheng et al., 2023). A slight tilt from left to right in Figure 2 and Figure 4 can be expected, as not all questions are accompanied by 10 options, with a considerable amount having less.
@@ 10.1 Scalability to other languages
As not every language has readily available datasets of even a few domains, we took an approach of using just translated datasets for all domains other than those used for localized and cultural knowledge addition. This would enable reusing the approach to build bilingual LLMs optimized for other languages as long as a proficient LLM supports the language to translate the texts fluently. Given that the performance of the models trained on a mix of real and translated data as well as just translated data are nearly identical as seen in Table 6 and Table 7. This technique could be scaled to other languages.
@@ 10.2 Model Efficiency
Unsloth’s version of phi-4 (Unsloth AI, 2023) with Llama architecture led to an improved performance but increased emissions. Our model resulted in lesser emissions during evaluation over the open-llm-leaderboard while improving the model’s performance. A comparison of our model to the original and unsloth’s phi-4 can be seen in Figure 3.
Figure 3: Emissions : open-llm-leaderboard evaluation
@@ Appendix A Model Replication
The hyper-parameters used for training can be seen below in Appendix B. The initial training attempts using a portion of the data (i.e 8% samples) were done on various different devices, All experiments combined including evals and training consumed an equivalent of 642 Hours on H200 SXM.
@@
Appendix B License
Our Qwen and Phi models are available through the same licenses as the models we used as a base, i.e., Apache-2.0 and MIT, respectively. The models can be accessed here.  [^4] Our Phi-4 model :https://huggingface.co/. The training datasets are publicly available here.  [^5] Datasets: https://huggingface.co/. Most datasets used for training have a copyleft license, with the rest having no license specified and being publicly available on Hugging Face.
[…]
@@ Appendix B License
[…]
@@
The initially collected dataset sources, sample sizes and the later used sample counts can be seen in Table 12 along with the ratios of each language. The sampling within each dataset is done at random using the seed specified in Appendix B. The samples were sorted in ascending order based on input size and the longest 600 samples in terms of input token count were added in the beginning of the training data.
Appendix C Datasets and Benchmarks Info
@@
Appendix D Results from other attempts
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Appendix E Model Choices